# Eclipsed X-ray Bursts from Magnetar SGR J1935+2154 and the Fireball Measurements

### v1

@@ Appendix A Data Reduction
Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM, X. Li et al., 2021; S. Xiao et al., 2022) is originally a constellation of two small X-ray and gamma-ray all-sky observatories funded by the Chinese Academy of Sciences. These two microsatellites, denoted as GECAM-A and GECAM-B, were launched in December 2020. Each GECAM satellite is equipped with 25 Gamma-ray detectors (GRDs) and 8 Charged Particle Detectors (CPDs). Each GRD operates in two readout channels: high gain (HG, 20–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The time resolution of GRDs is 0.1 $\mu\rm s$. We select detectors with detector-to-source angles of $\leqslant 50^{\circ}$ in our analysis. Given the dominated emission energy range of magnetar bursts ($\lesssim$100 keV), we focus on using the HG channel for data analysis. The GECAM data reduction tool is GECAMTools  [^2] https://github.com/zhangpeng-sci/GECAMTools-Public. Considering the lanthanum absorption edge at 38.89 keV for LaBr3, as this feature potentially leads to systematic uncertainties in spectra, we have ignored the data in the range of 35-40 keV (J.-J. He et al., 2023).
Fermi Gamma-ray Burst Monitor (GBM, C. Meegan et al., 2009) has a continuous broadband energy coverage which consists of 12 NaI detectors (8-1000 keV) and 2 Bismuth Germanate (BGO) detectors (0.2-40 MeV). GBM provides three kinds of science data: CSPEC with continuous high energy resolution (1024 ms and 128 energy channels), CTIME with continuous high time resolution (64 ms and 8 energy channels), and TTE (time tagged photon event) with temporal resolution 2 $\mu s$ and energy resolution 128 channels. We choose the detectors with good detector-to-source angles ($\leqslant 50^{\circ}$), and hence the highest signal-to-noise ratio A. C. Collazzi et al. (2015); L. Lin et al. (2020b). The data reduction tool used to generate light curves and spectral data is GBM Data Tools (A. Goldstein et al., 2022). Response files are generated by GBM Response Generator v1.13 for spectral analysis. Similar to GECAM, we excluded 30-40 keV data for NaI detectors to avoid the iodine K-edge at 33.17 keV (E. Bissaldi et al., 2009).
@@ Appendix B Spectral Analysis
The software elisa is employed for exploring and analyzing the spectral data (W.-C. Xue et al., 2024). In performing maximum likelihood fitting, the $\chi^{2}$-statistics are used as a likelihood value to be minimized, and spectral channels are grouped to ensure a minimum of one count per channel for the validity of the fit statistic. Furthermore, to evaluate the performance of different spectral models, we used two commonly adopted information criteria: the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC) (A. R. Liddle, 2007). Both criteria balance the goodness of fit and the complexity of the model, penalizing models with more free parameters. Lower AIC or BIC values indicate a more statistically favored model.
@@
where $k_{\mathrm{B}}T$ is the observed temperature in units of keV, $A_{\mathrm{obs}}$ is the observed emission area in km2 and $D_{10}$ is the distance to the source in units of 10 kpc. The distance of SGR J1935+2154 with $D\simeq 9\ {\rm kpc}$ is adopted (S.-Q. Zhong et al., 2020). This model involves the fewest assumptions. Our analysis shows that the BB temperature fluctuates slightly around a constant value. This temperature variation may result from statistical fluctuations caused by insufficient photon counts, or alternatively, it could stem from the dynamical fluctuations of the fireball. In any case, based on the assumption of a light-curve plateau, we further consider ignoring these fluctuations, directly assuming a constant temperature, and using this temperature to perform a joint fitting of time-resolved spectra during the plateau phase (i.e., the temperatures are tied (shared) across different time intervals). The logarithm of the likelihood for the joint fitting is defined as

### v2

@@ Appendix A Data Reduction
Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM, Li et al., 2021b; Xiao et al., 2022) is originally a constellation of two small X-ray and gamma-ray all-sky observatories funded by the Chinese Academy of Sciences. These two microsatellites, denoted as GECAM-A and GECAM-B, were launched in December 2020. Each GECAM satellite is equipped with 25 Gamma-ray detectors (GRDs) and 8 Charged Particle Detectors (CPDs). Each GRD operates in two readout channels: high gain (HG, 20–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The time resolution of GRDs is 0.1 $\mu\rm s$. We select detectors with detector-to-source angles of $\leqslant 50^{\circ}$ in our analysis. Given the dominated emission energy range of magnetar bursts ($\lesssim$100 keV), we focus on using the HG channel for data analysis. The GECAM data reduction tool is GECAMTools (zhang, 2024). Considering the lanthanum absorption edge at 38.89 keV for $\mathrm{LaBr_{3}}$, as this feature potentially leads to systematic uncertainties in spectra, we have ignored the data in the range of 35-40 keV (He et al., 2023).
Fermi Gamma-ray Burst Monitor (GBM, Meegan et al., 2009) has a continuous broadband energy coverage which consists of 12 NaI detectors (8-1000 keV) and 2 Bismuth Germanate (BGO) detectors (0.2-40 MeV). GBM provides three kinds of science data: CSPEC with continuous high energy resolution (1024 ms and 128 energy channels), CTIME with continuous high time resolution (64 ms and 8 energy channels), and TTE (time tagged photon event) with temporal resolution 2 $\mu s$ and energy resolution 128 channels. We choose the detectors with good detector-to-source angles ($\leqslant 50^{\circ}$), and hence the highest signal-to-noise ratio Collazzi et al. (2015); Lin et al. (2020b). The data reduction tool used to generate light curves and spectral data is GBM Data Tools (Goldstein et al., 2022). Response files are generated by GBM Response Generator v1.13 for spectral analysis. Similar to GECAM, we excluded 30-40 keV data for NaI detectors to avoid the iodine K-edge at 33.17 keV (Bissaldi et al., 2009).
@@ Appendix B Spectral Analysis
The software elisa is employed for exploring and analyzing the spectral data (Xue et al., 2025). In performing Bayesian fitting, the likelihood based on the $\chi^{2}$-statistic is used to compute the posterior distribution, and spectral channels are grouped to ensure a minimum of one count per channel. Furthermore, to evaluate the performance of different spectral models, we used two commonly adopted information criteria: the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC) (Liddle, 2007). Both criteria balance the goodness of fit and the complexity of the model, penalizing models with more free parameters. Lower AIC or BIC values indicate a more statistically favored model.
@@
where $k_{\mathrm{B}}T$ is the observed temperature in units of keV, $A_{\mathrm{obs}}$ is the observed emission area in km2 and $D_{10}$ is the distance to the source in units of 10 kpc. The distance of SGR J1935+2154 with $D\simeq 9\ {\rm kpc}$ is adopted (Zhong et al., 2020). This model involves the fewest assumptions. Our analysis shows that the BB temperature fluctuates slightly around a constant value. This temperature variation may result from statistical fluctuations caused by insufficient photon counts, or alternatively, it could stem from the dynamical fluctuations of the fireball. In any case, based on the assumption of a light-curve plateau, we further consider ignoring these fluctuations, directly assuming a constant temperature, and using this temperature to perform a joint fitting of time-resolved spectra during the plateau phase (i.e., the temperatures are tied (shared) across different time intervals). The logarithm of the likelihood for the joint fitting is defined as